# SpotVLM: Cloud-edge Collaborative Real-time VLM based on Context Transfer

### v1

@@ 4.3 Visual Focus
Grounding-Guided Visual Representation.
The ROIs returned by the LVLM guide the grounding of the SVLM on historical frames. Most VLMs, including the Qwen-2.5-VL-3B (Bai et al. 2025), use a Vision Transformer (ViT) that divides a resized input image into patches and encodes them into visual tokens (e.g., 1824 tokens for a 1600 × 900 image). However, many of these tokens are irrelevant to the reasoning task. To accelerate inference and enhance focus, ROIs are mapped to corresponding patches, and only task-relevant ones are selected. This selection is formally defined as:
@@
where $\mathcal{P}$ denotes the set of all image patches, and $p_{i}$ is a patch whose spatial overlap with the ROI is non-empty. In the case of multi-object recognition, the largest ROI is selected for each target category. For the case illustrated in the figure, the number of features selected for projection is reduced to 49, thereby reducing the computational cost.
nuScenes BDD100K MOT IPN Hand ACE
Micro-F1 Macro-F1 0-1 EM Micro-F1 Macro-F1 0-1 EM Acc. Acc.
SVLM Only (Local) 58.19 37.70 16.10 62.12 31.06 6.37 36.41 23.43
LVLM Only (Cloud) 45.76 32.04 20.75 49.65 25.85 20.28 23.91 61.14
Distributed VLM (Li et al. 2025b) 45.31 31.78 20.39 49.31 25.71 20.09 23.64 60.86
LAECIPS (Hu et al. 2024a) 59.84 39.64 24.29 62.47 32.14 21.05 39.40 48.71
ADAS (Hu et al. 2024b) 60.00 41.35 18.11 63.87 33.12 10.63 39.27 55.43
SpotVLM 69.96 48.19 27.32 71.19 36.26 23.88 45.11 62.75
Grounding-Tracked Visual Representation.
The ROIs provided by the LVLM support cross-frame tracking, guiding the SVLM to focus on relevant regions in subsequent frames. Due to limited grounding capability, lightweight SVLMs may struggle to localize objects even when their categories are known. However, given the temporal continuity of video, ROI features from frame T-d can be matched to visual features in frame T. Let $\mathcal{P}_{T-d}^{(c)}$ denotes the set of patch indices for category $c$ on frame T-d, and $\mathbf{v}_{i}^{(T-d)}$ is the visual feature of the $i$-th patch. Thus, feature vector of object category $c$ on frame T-d is:
@@
We compute the cosine similarity between ROI features from frame T-d and the visual features of frame T, thereby obtaining similarity maps $S^{(c)}$ for each potential category $c$ in frame T.
@@
where $v^{(T)}_{j}$ denotes the feature of the $j$-th patch on frame T, and $N$ is the total number of patches. As shown in Fig.3, the LVLM accurately detects and localizes small objects such as traffic cones and pedestrians in frame T-d, and their mapped similarity heatmaps on frame T effectively indicate the potential locations after viewpoint shift, providing precise visual guidance for the SVLM. This cross-frame computation yields patch-level similarity scores can be aggregated into a unified similarity vector $s$, which can be defined as:
@@
Then $s$ can be used to construct the grounding-tracked visual weight $W_{s}$, formulated as follows:
@@
where $\mu$ and $b$ are tunable scaling parameters used to control the visual weights.
@@
$t\leftarrow(1-\alpha)\cdot t+\alpha\cdot W_{s}\cdot t.$ (10)
@@ 5.1 Experiments Setting
We evaluate the proposed architecture on three real-time tasks.
@@
Real-time multi-object recognition. (i) nuScenes (Caesar et al. 2019): We sample camera images at 1 FPS from this autonomous driving dataset. The processed data includes 150 scenes (20 seconds each) with 10 object categories. (ii) BDD100K (MOT) (Yu et al. 2020): We sample the driving video frame at 1 FPS. The processed data includes 200 scenes (40 seconds each) with 10 object categories.
@@
Real-time gesture recognition. IPN Hands (Benitez-Garcia et al. 2021): A gesture recognition dataset with 14 gesture classes, sampled at 2 FPS from 52 close-range videos (avg. 120s each) to capture dynamic motion.
@@
Real-time video frame caption. Actions for Cooking Eggs (ACE) (Shimada et al. 2012): A kitchen scene dataset with 8 cooking actions, sampled at 1 FPS from 10 videos (avg. 194s each).
Metrics.
@@
We adopt commonly used evaluation metrics (Durand, Mehrasa, and Mori 2019; Nam et al. 2017; Huang et al. 2017) for multi-object recognition task, which are Micro-F1, Macro-F1 and 0-1 Exact Match (0-1 EM). Specifically, Micro-F1 reflects overall performance across all instances, and Macro-F1 captures class-wise performance equally. Additionally, 0-1 EM metric equals 1 only if the predicted result exactly matches the ground truth label for the current frame.
@@
Following 0-1 EM, we model both gesture recognition and video frame captioning as frame-level single-label classification tasks, and use accuracy (Acc) as the evaluation metric.
Note that the evaluated tasks operate under real-time constraints (1–2 FPS). If the system fails to return a valid result within one second, the output for that frame is considered empty.
@@
Implementation Details.
In the Visual Focus Module, the parameters $\mu$, $b$ and $\alpha$ are set to 0.5, 0.75, and 0.1, respectively. To balance processing speed and model performance, we retain two previous frames as context. SpotVLM consistently treats the cloud feedback as the farthest frame in the historical context of SVLM.
In addition, to evaluate the effectiveness of different collaborative architectures, we use the official pretrained versions of both the LVLMs and SVLMs without any additional fine-tuning. During inference, the outputs and response latencies of the large model are recorded for each frame to ensure fair comparison of collaboration strategies. All experiments on the edge side are conducted using a GeForce RTX 4090 GPU with 24 GB of memory.
CRM GGVR GTVR Micro-F1 Macro-F1 0-1 EM
66.87 43.07 24.39
71.46 48.64 29.98
72.15 50.85 30.93
73.02 51.56 33.17
@@
SVLM LVLM Micro-F1 Macro-F1 0-1 EM
Baseline SpotVLM Baseline SpotVLM Baseline SpotVLM
Qwen2.5-VL-3B Qwen2.5-VL-72B 66.87 73.02 43.07 51.56 24.39 33.17
MiniCPM-o-2.6-int4 Qwen2.5-VL-72B 67.68 73.56 43.47 49.63 31.38 35.71
Qwen2.5-VL-3B GPT-4o 66.34 74.25 47.85 53.74 22.53 30.79
Table 3: Adaptability Study. Testing the SpotVLM architecture with various SVLM–LVLM combinations on nuScenes.
5.2 Overall Performance
To evaluate the effectiveness of different architectures, we assess six methods under a Wi-Fi network environment: locally deployed Qwen2.5‑VL‑3B (Bai et al. 2025), cloud-based Qwen2.5‑VL‑72B (Bai et al. 2025), Distributed VLM, LAECIPS, ADAS, and SpotVLM. The results are presented in Tab.1.
Due to cloud latency, methods that rely entirely on large models (e.g., Qwen2.5-VL-72B, Distributed VLM) yield lower F1 scores on real-time tasks than collaborative approaches with small models. However, the 0–1 EM of Qwen2.5-VL-72B remains higher than that of the small model alone, showing its superior accuracy when responses are timely. Additionally, Qwen2.5-VL-3B produces visual features larger than the original image, making Distributed VLM, which transmits features to the cloud, slightly less effective than running full inference on the cloud model.
Collaborative architectures like LAECIPS and SpotVLM perform consistently well across all four datasets. When cloud responses are timely, they effectively leverage large-model outputs, achieving better overall metrics than Qwen2.5-VL-3B.
ADAS adopts a latency-adaptive offloading strategy. In the first three datasets, long outputs from multi-object recognition and dual-image inputs in gesture recognition result in high cloud latency. Under these conditions, ADAS relies more on the small model, yielding higher F1 scores but lower 0–1 EM compared to LAECIPS. In the fourth dataset, shorter outputs and fewer images per inference lead to more timely cloud responses. ADAS thus relies more on the output of the LVLM. Meanwhile, LAECIPS depends more on the SVLM for simple samples in this dataset, resulting in slightly lower performance than ADAS.
To ensure a fair comparison across methods, we recorded the number of cloud service invocations for ADAS and LAECIPS, and used their average values to set the interval for SpotVLM. Benefiting from its ability to effectively leverage the high-quality outputs of the LVLM under latency constraints, SpotVLM consistently achieved the best results across all tasks and collaborative strategies.
To provide a more intuitive comparison, we select one representative example from each of the four datasets for qualitative analysis. As shown in the Fig.4, for the real-time multi-object recognition task, all three methods rely on the outputs of the SVLM, with SpotVLM achieving the best performance due to guidance from the historical results of the large model. In the gesture recognition task, as two images need to be transmitted and processed simultaneously, the cloud-based large model fails to return valid results in real time, leading to No response for ADAS. In contrast, SpotVLM leverages the history-enhanced small model to produce accurate classification results. For the video-frame captioning task, both ADAS and SpotVLM adopt the results quickly returned by the large model.
5.3 Ablation Study
We conduct a progressive ablation study on the nuScenes dataset to evaluate the individual and combined contributions of Contextual Replacement Module (CRM), Grounding-Guided Visual Representation (GGVR), and Grounding-Tracked Visual Representation (GTVR) in the SpotVLM framework. To isolate component effects, all image frames are uploaded, eliminating the influence of frame selection and upload strategies.
As shown in the Tab.2, the first row corresponds to the baseline where the large and small models operate in parallel—using the LVLM output when it is returned in time, and otherwise defaulting to the SVLM. This baseline represents the upper-bound performance of current cloud–edge collaborative VLM architectures. Compared to this baseline, we incrementally integrate each proposed module. Results show that CRM yields the most significant performance gain. GGVR is used to reduce the number of visual tokens, while GTVR guides the SVLM to maintain consistent detection of object categories present in the historical context. Building on CRM, the integration of these two methods further enhances overall system performance across all three metrics.
5.4 Adaptability Study
@@
To evaluate the effectiveness of the proposed framework under different combinations of model capacities, we conducted additional experiments by replacing both the small and large models. As in the ablation study, all image frames are uploaded and processed by the large model, and the upper bound is represented by a parallel inference strategy that always adopts the output of large model when available. Specifically, we adopt MiniCPM-o (Yao et al. 2024) as the SVLM on the edge side and GPT-4o (Hurst et al. 2024) as the LVLM on the cloud side. As shown in the Tab.3, the proposed SpotVLM coordination strategy consistently improves real-time inference performance across model configurations. With the continuous advancement of small and large VLMs, more powerful models can be seamlessly integrated into the SpotVLM framework to further enhance system capabilities.
Figure 5: Latency simulation. Manual simulation of total processing latency on the cloud LVLM. Red pentagrams denote the performance bounds of LVLM and SVLM; the dashed line shows the SVLM-only performance.
5.5 Latency Analysis
To evaluate the impact of cloud latency on SpotVLM, we conduct experiments on the nuScenes dataset using Qwen-2.5-VL-3B as the edge-side SVLM and preserving the outputs of Qwen-2.5-VL-72B as the cloud-side LVLM. We manually simulate various latency conditions by setting different return delays for the LVLM responses.
When the latency is 1 second, which means that the LVLM output always arrives within the real-time threshold, the system achieves performance equivalent to directly using the LVLM (denoted by the star marker). Starting from a 2-second delay, the LVLM result can only be incorporated as historical context for guiding the SVLM. Even under this condition, SpotVLM significantly outperforms the standalone SVLM (indicated by the dashed line). As latency increases, the temporal correlation between frames weakens, leading to a gradual decline in overall performance. However, SpotVLM continues to maintain a notable advantage even under a 9-second delay.
This experiment reveals that the performance upper bound is achieved when the LVLM returns results in real time for every frame, while the lower bound corresponds to using only the SVLM. The proposed collaborative architecture offers strong robustness under fluctuating latency conditions. Moreover, the results also suggest that cloud-side LVLM invocation can be dynamically scheduled. By adjusting the upload interval according to task requirements and performance targets, the system can reduce the frequency of cloud model calls without significantly compromising overall effectiveness.

### v2

@@ 3.3 Visual Focus
Figure 3: ROI-based region retrieving via cosine similarity. Given ROIs of traffic cone and pedestrian from frame T – d, cosine similarity highlights related regions in frame T to guide the attention of the SVLM.
3.3.1 Grounding-Guided Visual Representation
The ROIs returned by the LVLM guide the grounding of the SVLM on historical frames. Most VLMs, including the Qwen-2.5-VL-3B [Qwen2.5-VL], use a Vision Transformer (ViT) [dosovitskiy2020image] that divides a resized input image into patches and encodes them into visual features (e.g., 1824 features for a 1600 × 900 image). However, many of these features are irrelevant to the reasoning task. To accelerate inference and enhance focus, ROIs are mapped to corresponding patches, and only task-relevant ones are selected. This selection is formally defined as:
@@ 3.3.1 Grounding-Guided Visual Representation
where $\mathcal{P}$ denotes the set of all image patches, and $p_{i}$ is a patch whose spatial overlap with the ROI is non-empty. In the case of multi-object recognition, the largest ROI is selected for each target category. For the case illustrated in Fig. 2, the number of visual tokens selected after projection is reduced to 49, thereby reducing the computational cost.
3.3.2 Grounding-Retrieved Visual Representation
The ROIs provided by the LVLM support cross-frame retrieving, guiding the SVLM to focus on relevant regions in subsequent frames. Given the temporal continuity, ROI features from frame T - d can be matched to visual features in frame T. Let $\mathcal{P}_{T-d}^{(c)}$ denotes the set of patch indices for category $c$ on frame T - d, and $\mathbf{v}_{i}^{(T-d)}$ is the visual feature of the $i$-th patch. Thus, feature vector of object category $c$ on frame T-d is:
@@ 3.3.2 Grounding-Retrieved Visual Representation
We compute the cosine similarity between ROI features from frame T - d and all visual features from frame T, thereby obtaining similarity maps $S^{(c)}$ for each potential category $c$ in frame T.
@@
where $v^{(T)}_{j}$ denotes the feature of the $j$-th patch on frame T, and $N$ is the total number of patches. As shown in Fig.3, the LVLM accurately detects and localizes small objects such as traffic cones and pedestrians in frame T - d, and their mapped similarity heatmaps on frame T effectively indicate the potential locations after viewpoint shift, providing precise visual guidance for the SVLM. This cross-frame computation yields patch-level similarity scores, which can be aggregated into a unified similarity vector $s$, defined as:
@@
Then $s$ can be used to construct the grounding-retrieved saliency weight $W_{s}$, formulated as follows:
@@
where $\mu$ and $b$ are tunable scaling parameters used to control the visual weights. Here, we enforce $0.5\mu+b=1$ in our experiments.
@@
$t\leftarrow\alpha\cdot t+(1-\alpha)\cdot W_{s}\cdot t.$ (11)
@@
3.4 Keyframe Decision
Figure 4: Experimental latency evaluation of cloud services over a 5G network.
Under a 5G network environment, we invoke the Qwen-VL-2.5-72B API hosted on Alibaba Cloud and record the per-frame response latency, as illustrated in the Fig.4. The results show that adjacent frames exhibit similar latency patterns. To enhance the efficiency of LVLM invocations and dynamically assess network conditions, the system continuously monitors how many preceding frames have not yet received responses. Based on this observation, we introduce a hyperparameter $K_{delay}$, which serves as a threshold: when the number of unreturned frames accumulates to $K_{delay}$, the system infers that the network condition has deteriorated and skips uploading the current frame.
We further regulate LVLM call frequency with a hyperparameter $K_{call}$. For frame $T$, if the most recent LVLM result is from frame TiT-i with $i\leq K_{call}$, that history is considered sufficiently fresh to guide the SVLM, and frame $T$ is not uploaded. This rule takes precedence over $K_{delay}$.
@@ 4.1.1 Datasets
We evaluate the proposed architecture on three real-time tasks across four datasets.
@@
Real-time multi-object recognition. (i) nuScenes [nuscenes2019]: We sample images at 1 FPS from this autonomous driving dataset, which finally includes 156 scenes (avg. 20s each) with 10 object categories. (ii) BDD100K (MOT) [yu2020bdd100k]: We sample the driving video frame at 1 FPS, which includes 200 scenes (avg. 40s each) with 10 object categories.
@@
Real-time gesture recognition. IPN Hands [bega2020IPNhand]: A gesture recognition dataset with 14 gesture classes, sampled at 2 FPS from 52 close-range videos (avg. 120s each) to capture dynamic motion.
@@
Real-time video frame caption. Actions for Cooking Eggs (ACE) [shimada2012kitchen]: A kitchen scene dataset with 8 cooking actions, sampled at 1 FPS from 10 videos (avg. 194s each).
4.1.2 Metrics
@@
We adopt commonly used evaluation metrics [durand2019learning, nam2017maximizing, huang2017multi] for multi-object recognition task, which are Micro-F1, Macro-F1 and 0-1 Exact Match (0-1 EM). 0-1 EM equals 1 only if the predicted result exactly matches the ground truth label.
@@
We model both gesture recognition and video frame captioning as frame-level single-label classification tasks, and use accuracy (Acc) as the evaluation metric.
Note that the evaluated tasks operate under real-time constraints (1–2 FPS). Following DriveVLM [tian2024drivevlm], we set $\tau=1$ in experiments.
4.1.3 Implementation Details
Experiments are conducted over a commercial 5G network. The average uplink bandwidth was 211 Mbps, downlink 72 Mbps, with measured round-trip latency around 42 ms. To remove variability from network jitter and queuing, we execute the cloud-side LVLM on every frame and log its streaming outputs and the per-frame latency, as shown in Fig. 4. All collaboration strategies are evaluated against these identical per-frame traces for a fair comparison.
To balance processing speed and model performance, we retain two previous frames as context. All experiments on the edge side are conducted using a GeForce RTX 4090 GPU with 24 GB of memory.
The following abbreviations are used in this section: Grounding-Guided Visual Representation (GGVR), and Grounding-Retrieved Visual Representation (GRVR).
4.2 Hyperparameter Study
LVLM Call. To balance the system performance and the cost of invoking the LVLM, we introduce two hyperparameters, $K_{delay}$ and $K_{call}$, in the Keyframe Decision module. A larger $K_{delay}$ leads to more frequent LVLM invocations. We conduct evaluations on the nuScenes dataset, and the resulting system performance is shown in Fig.5(a). $K_{delay}=4$ is selected as a trade-off choice. Under the 5G network, we recorded the LVLM utilization under different $K_{call}$ settings, as shown in Tab.1. LCR (LVLM Call Ratio) is the fraction of inference steps that trigger the cloud LVLM. LHL (LVLM History Lag) is the average frame offset $i$ between the current frame $t$ and the reused LVLM reply from frame tit-i, computed over all frames that actually consume LVLM history. When $K_{call}$ increases from 2 to 3, the LVLM invocation frequency drops sharply, while the overall system performance remains relatively stable as shown in Fig.5(b).
Visual Guidance. In the GRVR module, $\alpha$ and $\mu$ are introduced to regulate the visual guidance. Considering multi-metric performance and robustness, we adopt $(\alpha,\mu)=(0.7,1.0)$.
Figure 5: Absolute performance gain of edgeVLM over SVLM across different settings. (a) shows the impact of $K_{delay}$ with $K_{call}=0$. (b) represents the impact of $K_{call}$ with $K_{delay}=4$.
$K_{call}$ 1 2 3 4 5 6 7 8 9
LCR(%) 69.1 38.9 27.3 23.5 19.0 18.4 17.8 14.4 13.0
LHL(s) 1.89 2.56 3.31 3.79 4.12 4.54 5.14 5.48 5.66
Table 1: Impact of $K_{call}$ on LVLM utilization.
Method nuScenes BDD100K MOT IPN Hand ACE
Micro-F1 Macro-F1 0-1 EM LCR Micro-F1 Macro-F1 0-1 EM LCR Acc LCR Acc LCR
SVLM Only (Edge) 55.36 38.51 13.80 0 61.56 30.79 6.59 0 36.68 0 22.86 0
LVLM Only (Cloud) 43.84 33.81 18.98 100 49.79 24.59 21.77 100 22.55 100 60.14 100
Distributed VLM [li2025distributed] 43.35 33.63 18.84 100 48.28 23.55 20.89 100 21.47 100 59.43 100
LVLM Call Ratio (20-30%)
LAECIPS [hu2024laecips] 58.06 39.01 17.04 29.3 61.70 30.48 12.88 28.7 38.86 28.8 32.57 28.0
ADAS [hu2024cloud] 56.43 40.28 18.43 28.2 62.10 31.54 13.61 29.5 37.09 25.6 33.29 28.5
edgeVLM 64.49 48.00 21.99 27.3 68.44 34.76 23.66 28.4 46.20 20.7 34.71 25.9
LVLM Call Ratio (35-45%)
LAECIPS 57.24 39.39 19.58 44.5 61.88 30.90 15.98 44.6 37.23 44.0 40.29 44.7
ADAS 57.52 41.13 20.03 43.7 62.55 30.62 16.88 44.2 35.87 43.8 39.71 43.6
edgeVLM 66.22 50.04 24.51 43.2 70.63 36.41 26.88 43.7 47.42 35.4 38.86 35.1
LVLM Call Ratio (50-60%)
LAECIPS 56.28 39.06 21.18 55.3 61.91 30.28 17.56 59.2 34.51 58.7 45.71 54.0
ADAS 58.03 42.23 22.13 55.9 62.88 30.62 19.99 57.4 35.33 56.2 43.23 54.8
edgeVLM 67.24 51.35 25.00 52.8 71.70 37.31 29.16 57.3 49.73 55.7 46.14 52.9
LVLM Call Ratio (100%)
edgeVLM 71.15 57.11 29.41 100 74.12 39.66 35.08 100 52.45 100 63.00 100
@@ 4.2 Hyperparameter Study
4.3 System-level Evaluation
4.3.1 Overall Performance
To evaluate the effectiveness of different architectures, we assess six methods: edge-based Qwen2.5‑VL‑3B, cloud-based Qwen2.5‑VL‑72B [Qwen2.5-VL], Distributed VLM, LAECIPS, ADAS, and edgeVLM. The results are presented in Tab.2.
For the fourth task, single-image inputs and short outputs allow rapid LVLM feedback, making full LVLM invocation clearly advantageous. For the other tasks, due to cloud latency, methods that rely entirely on LVLM yield lower F1 scores than collaborative approaches. However, the 0–1 EM of LVLM Only remains higher than that of SVLM Only, showing its superior accuracy when responses are timely. Additionally, Qwen2.5-VL-3B produces visual features larger than the original image, making Distributed VLM less effective than directly invoking the LVLM.
Figure 7: Adaptability Study. Testing the edgeVLM architecture with various SVLM–LVLM combinations on nuScenes. VLM-R1 is a small model further enhanced by reinforcement learning and knowledge distillation.
By tuning the hyperparameters, we evaluate the system performance of edgeVLM under different LCR settings and compare it with alternative architectures. ADAS adopts latency-adaptive offloading strategies, where LVLM invocation depends on the current cloud service state. As a result, it tends to use a relatively low LCR on the first three datasets under realistic deployment. In contrast, LAECIPS triggers LVLM calls based on task difficulty and thus chooses a higher LCR on these datasets. Overall, however, both schemes deliver lower practical performance than edgeVLM, which more effectively exploits historical information even with a comparatively low LCR.
To provide a more intuitive comparison, we select four representative examples for qualitative analysis. As shown in Fig.6, for the real-time multi-object recognition task, all three methods rely on the SVLM, with edgeVLM achieving the best performance due to guidance from the historical results of the LVLM. In the gesture recognition task, as two images need to be transmitted and processed simultaneously, the LVLM fails to return valid results in real-time, leading to No response for ADAS. For the video-frame captioning task, both ADAS and edgeVLM adopt the results quickly returned by the LVLM.
4.3.2 Latency Analysis
System Stage Time (ms)
LVLM Only Total 2110.83
SVLM Only Total 581.71
edgeVLM CRM-induced Overhead 19.93
VFM-induced Overhead 31.27
Inference 511.20
Total 562.40
Table 3: Execution time breakdown by system and stage (ms).
To assess module-induced computational overhead, we report the average execution time of different systems in Tab. 3. CRM and VFM introduce additional processing latency. However, by reducing the number of visual tokens, VFM lowers the compute over the entire inference pipeline and offsets part of this added overhead. As a result, edgeVLM attains a lower total latency than running the SVLM. Moreover, efficiency details are presented in Tab.4.
System GPU Memory (MiB) Prefill Time (ms) Time / Token (ms)
SVLM Only 10411 169.38 46.86
edgeVLM 9589 166.81 44.75
Table 4: Efficiency analysis of edgeVLM.
4.4 Method-level Evaluation
4.4.1 Ablation Study
We conduct a progressive ablation study on the nuScenes dataset to evaluate the individual and combined contributions of different modules. To isolate component effects, all image frames are uploaded, eliminating the influence of frame selection and upload strategies.
As shown in the Tab.5, the first row corresponds to the baseline where the LVLM and SVLM operate in parallel—using the LVLM output when it is returned in time, and otherwise defaulting to the SVLM. This baseline represents the upper-bound performance of existing cloud–edge collaborative VLM architectures. Compared to this baseline, we incrementally integrate each proposed module. Results show that CRM yields the most significant performance gain. Building on CRM, the integration of the other two methods further enhances overall system performance across all three metrics.
CRM GGVR GRVR Micro-F1 Macro-F1 0-1 EM
64.36 46.55 22.35
70.47 55.43 28.85
70.77 55.58 29.20
71.15 57.11 29.41
@@ 4.4.1 Ablation Study
4.4.2 Adaptability Study
@@
To evaluate the effectiveness of the proposed framework under different combinations of model capacities, we conduct additional experiments by replacing both the SVLMs and LVLMs. All image frames are uploaded as described in Sec.4.4.1. Specifically, we add MiniCPM-o [yao2024minicpm] and VLM-R1 [shen2025vlm] as the SVLMs and GPT-4o [hurst2024gpt] as the LVLM. As shown in the Fig.7, the proposed edgeVLM coordination strategy consistently improves real-time inference performance across model configurations. VLM-R1 is an SVLM enhanced through distillation and reinforcement learning. Experiments demonstrate that powerful models can be seamlessly integrated into the edgeVLM framework to further enhance system capabilities.